OpenStudy (anonymous):

4sin2x-1 = 1 find all possible solutions in degrees and radians from 0 - 2pi

OpenStudy (mertsj):

Is it sin 2x or sin^2 x?

OpenStudy (anonymous):

4 sin (2x) -1 = 1

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

4sin2x-1=1 4sin2x=2 sin2x=2/4 sin2x=1/2 solve from there

OpenStudy (anonymous):

there are four solutions?!?

OpenStudy (anonymous):

those are the steps

OpenStudy (anonymous):

sorry i would rather do it the way my eacher taught me. so i hade sin 2x = 1/2. now what?

OpenStudy (anonymous):

the answers are pi/6, 11pi/6, 5pi/6

OpenStudy (anonymous):

if i plug in sin 1/2 into my calculator i get 30 degrees then divide by 2 to get 15 degrees as reference angle?

OpenStudy (mertsj):

\[\sin(2x)=\frac{1}{2}\]\[2x=\frac{pi}{6}, \frac{5pi}{6}, \frac{13pi}{6}, \frac{17pi}{6}\] \[x=\frac{pi}{12}, \frac{5pi}{12}, \frac{13pi}{12}, \frac{17pi}{12}\]

OpenStudy (anonymous):

ugh guys shouldn't there only be 2 solutions because sin is positive so it must be in quadrant 1 and 2!!!!

OpenStudy (anonymous):

you don't understand this has confused me for months and my teacher doesn't care

OpenStudy (anonymous):

yes sorry my mistake pi/6 and 5pi/6 are your answers

OpenStudy (mertsj):

No. Since the argument is 2x that means you have to make two complete revolutions. You will be dividing by 2 so to get all the answers between 0 and 2 pi after dividing by 2, you have to start out with all the answers between 0 and 4 pi

OpenStudy (mertsj):

@Brent0423 Are you sure the information you are disseminating is correct?

OpenStudy (anonymous):

it says between 0-2pi

OpenStudy (mertsj):

All 4 of those answers are between 0 and 2 pi

OpenStudy (anonymous):

omg im so screwed

OpenStudy (mertsj):

Why?

OpenStudy (anonymous):

i really don't get this. its supposed to have 2 solutions

OpenStudy (mertsj):

Who said?

OpenStudy (anonymous):

because sin is positive

OpenStudy (anonymous):

so Q1 Q2

OpenStudy (mertsj):

Plug those 4 answers I gave you into your calculator and see if they make the original equation true.

OpenStudy (anonymous):

how..?

OpenStudy (anonymous):

wait wouldn't it be 15 degree not 30?

OpenStudy (mertsj):